OpenStudy (thomas5267):

Too tired to think. Will give medal to correct answer. This is not a homework question. Consider two polynomial p(x) and d(x). The degree of p(x) is larger than d(x). Divide p(x) with d(x) and let the quotient be q(x) and remainder be r(x). Find a relationship between degree of p(x), q(x), d(x) and r(x).

OpenStudy (thomas5267):

Wait. I screwed that up.

ganeshie8 (ganeshie8):

okie

OpenStudy (thomas5267):

Put it in simple wordings, find a relationship between the degree the dividend, the divisor, the quotient and the remainder.

OpenStudy (thomas5267):

@ganeshie8 The question is fixed.

ganeshie8 (ganeshie8):

p(x) = q(x)*d(x) + r(x)

ganeshie8 (ganeshie8):

you want to get a relationship between their degrees is it

OpenStudy (thomas5267):

Yes.... I feel dreadful because of sleep deprivation ahhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh.

ganeshie8 (ganeshie8):

degree of q = p-d

ganeshie8 (ganeshie8):

that much is clear right ?

OpenStudy (thomas5267):

Yes. Then deg(q)+deg(d)=deg(p)? I am tirrrrrrrrrrrrrrrrrrrrrrrrrrrrrrrred.

ganeshie8 (ganeshie8):

yes that looks good!

ganeshie8 (ganeshie8):

And, deg(r) < deg(d)

OpenStudy (thomas5267):

THIS IS NOT WORKING AS INTENDED ARGHHHHHHHHHHH! I thought figuring this out could help me distinguish which part of the result of the synthetic division is the quotient and which part of the result is the remainder. How do I figure that out?

ganeshie8 (ganeshie8):

are you dividing a polynomial by a linear polunomial like `ax+b` ?

OpenStudy (thomas5267):

Or a irreducible quadratic over the reals.

ganeshie8 (ganeshie8):

Consider below division : \(\large \dfrac{x^3+2x+1}{x^2+1}\) x^3+2x+1 = `x` * `(x^2+1)` + `x+1`

ganeshie8 (ganeshie8):

ohk...

ganeshie8 (ganeshie8):

everything after first "p-d" elements is a remainder

ganeshie8 (ganeshie8):

can u copy paste ur last row of synthetic division ?

OpenStudy (thomas5267):

Wait. Isn't everything after the first (p-d+1) element is the remainder?

OpenStudy (thomas5267):

It took me seven seconds to calculate 3-1+1, which is slow. I suppose I am just tired?

ganeshie8 (ganeshie8):

Ah yes (p-d+1) looks correct because of the constant term !

ganeshie8 (ganeshie8):

lol happens haha! you must have tried something like 3-1+1 = 2+1-1+1 = 2+0+1 = 3 ?

OpenStudy (thomas5267):

I do it like this: 3-1+1 =2+1 (2 seconds) =3 (2 seconds) Then I have to do it again because I usually make stupid mistakes! That took me around 8 seconds total actually!

OpenStudy (thomas5267):

If you want a laugh, this is a good article on science and evolution. http://sciencealert.com.au/features/20141509-26177.html

ganeshie8 (ganeshie8):

lol that's a convincing theory :P ty for shairng ;)